A biologist collected 201 fern and moss samples. There were 85 more moss than fern samples. How many fern samples did the biologist collect?
Answer by lawdej(58)

let fern be f and moss be m
m+f=201 and m=f+85
f+85+f=201
2f=201-85
2f=116
f=116/2
f=58
so the biologist was collected 58 fern samples.